What is the equation, in slope intercept form, of the line that is parallel to the line x-3y=9 and passes through the point (6,3)?
Answer by mananth(16946)

1 x -3 y = 9
Find the slope of this line
make y the subject
-3 y = -1 x + 9
Divide by -3
y = 1/ 3 x -3
Compare this equation with y=mx+b
slope m = 1/ 3 0.8
The slope of a line parallel to the above line will be the same
The slope of the required line will be 1/ 3
m= 1/ 3 ,point ( 6 , 3 )
Find b by plugging the values of m & the point in y=mx+b

3 = 2 + b
b= 1
m= 1/3
Plug value of the slope and b in y = mx +b
The required equation is y = 1/3 x + 1

m.ananth@hotmail.ca